A question in Probability & Statics:-


At a checkout counter, customers arrive at an average of 1.5

per minute. Assuming poisson distribution

1- The probability of no arrival in two minutes is.....

2- The variance of the number of arrivals in two minutes

is.....

3- Suppose X has binomial distribution with n=1000 and p

=0.002 , Then P(x=3)=....


1
Expert's answer
2022-06-02T14:53:22-0400

1. "X\sim Po(\lambda t), \lambda t=1.5(2)=3"


"P(X=0)=\dfrac{e^{-3}(3)^0}{0!}=0.049787"

2.


"Var(X)=\lambda t=3"

3. When the value of "n"  in a binomial distribution is large and the value of "p" is very small, the binomial distribution can be approximated by a Poisson distribution. If "n > 20" and "np < 5" or "nq < 5" then the Poisson is a good approximation.

Check


"n=1000>20, np=1000(0.002)=2<10"


Then "\lambda=np=2"

"X\sim Po(2)"

"P(X=3)=\dfrac{e^{-2}(2)^3}{3!}""=0.180447"